# Explainable Procedural Mistake Detection

### v1

@@ Abstract
Automated task guidance has recently attracted attention from the AI research community. Procedural mistake detection (PMD) is a challenging sub-problem of classifying whether a human user (observed through egocentric video) has successfully executed the task at hand (specified by a procedural text). Despite significant efforts in building resources and models for PMD, machine performance remains nonviable, and the reasoning processes underlying this performance are opaque. As such, we recast PMD to an explanatory self-dialog of questions and answers, which serve as evidence for a decision. As this reformulation enables an unprecedented transparency, we leverage a fine-tuned natural language inference (NLI) model to formulate two automated coherence metrics for generated explanations. Our results show that while open-source VLMs struggle with this task off-the-shelf, their accuracy, coherence, and dialog efficiency can be vastly improved by incorporating these coherence metrics into common inference and fine-tuning methods. Furthermore, our multi-faceted metrics can visualize common outcomes at a glance, highlighting areas for improvement.
Explainable Procedural Mistake Detection
@@
Figure 1: Overview of self-reflective explanatory dialog between a foundational vision-and-language model (VLM) and itself to gather visual information and facilitate explainable procedural mistake detection (PMD).
@@ 1 Introduction
The problem of automated, interactive task guidance has recently attracted attention in the AI research community Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024); Bohus et al. (2024), stemming from significant efforts to build and learn from large-scale procedural video datasets Zhou et al. (2018); Damen et al. (2018); Miech et al. (2019); Grauman et al. (2022). A successful task guidance agent can observe a human user through video and interact with them through language to guide them through completing a task. One key component of such an agent is procedural mistake detection (PMD): the ability to detect when the user’s actions deviate from a procedural text, e.g., a recipe or instruction manual. To achieve this, a system must apply physical commonsense reasoning to anticipate the success conditions for actions in this procedural text, then verify them by extracting relevant physical state information from the visual scene.
@@
Prior work in PMD has explored two approaches: one thread of work has fine-tuned primarily vision-based classifiers without incorporating language Wang et al. (2023a); Peddi et al. (2024), while another has applied foundational language models (LMs) and vision-and-language models (VLMs) to this problem Du et al. (2023); Bao et al. (2023). However, both types of approaches have failed to achieve a viable level of accuracy in detecting mistakes. Qualitatively, Bao et al. (2023) find that while the web-scale multimodal pre-training of foundational VLMs enables coverage of a wide variety of procedures, they often produce noisy, vague, or otherwise insufficient information from visual scenes to facilitate PMD. This capability to extract and reason over key task-relevant visual information is crucial to PMD, but prior work has largely overlooked it, instead targeting binary or categorical classification tasks in their system design and quantitative evaluations. Consequently, the reasons for VLMs’ decisions are opaque, hindering practical use [^1] Since VLMs are generally unreliable procedural mistake detectors, transparency is crucial to ensure the human user can understand the system’s reasoning process and act on, disregard, or even correct its outputs accordingly. and the identification of areas for improvement.
@@
To promote the development of PMD systems with greater transparency and coherence, we propose the reformulated problem of explainable PMD: given a procedural text and egocentric video frame, VLMs must not only decide whether a mistake has occurred, but additionally explain their decision with visual information from the scene, specifically through an iterative self-dialog of generated questions and answers which condition their decision. [^2] Of the many possible forms this explanation could take, we choose this one because of its resemblance to visual question answering (VQA; Antol et al., 2015) and visual dialog Das et al. (2017), tasks that have long been studied in the NLP and computer vision communities, and recent open-source VLMs are often trained and perform well on Dai et al. (2023a); Liu et al. (2023); Dubey et al. (2024). We define two automated metrics for the coherence of generated explanations based on a fine-tuned natural language inference (NLI) model. To lay the groundwork for research into coherent explainable PMD, we apply three natural interventions to off-the-shelf VLMs: (1) we use our metrics to re-rank candidate questions generated by VLMs, (2) we harness their in-context learning capability by prompting them with human-written examples to augment the space of candidate questions, and (3) we fine-tune VLMs to generate more coherent questions using our metrics. Our results show that these interventions can significantly improve VLMs’ accuracy, coherence, and even the explanation efficiency. We lastly show how our multi-faceted metrics for this problem reveal previously impossible insights into VLMs’ reasoning processes, enabling possible targeted improvement in downstream engineering of task guidance systems.
@@ 2 Problem Formulation and Dataset
In this section, we define the extended problem of explainable PMD in an approachable manner for VLMs, then we introduce a benchmark dataset we curated for evaluating explainable PMD.
2.1 Defining Explainable PMD
Formally, the inputs for PMD are a short procedural text $T$ and a single video frame $F$, which may or may not visualize the successful completion of the procedure described in $T$. Given these inputs, a system should return a binary mistake detection decision $y$ for whether the procedure has been successfully completed ($y=0$ indicates success, and $y=1$ indicates a mistake). In explainable PMD, it must additionally generate an explanation for the decision $E=(\mathcal{Q},\mathcal{A})$, where $\mathcal{Q}=\{Q_{1},Q_{2},\dots,Q_{n}\}$ and $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$ are respectively sequences of $n$ yes-no questions and their predicted yes-no answers.
As shown in Figure 1, we instantiate this problem through an iterative self-reflective dialog of questions and answers. This dialog consists of several steps: proposing questions to ask about the procedure at hand (visual question generation or VQG), answering them (visual question answering or VQA), and making a success/mistake classification for whether the procedure has been completed in the given video frame. This structure goes beyond past approaches for PMD using VLMs; while Du et al. (2023) only elicited success classification, Bao et al. (2023) used procedure-agnostic prompts to caption images before making a classification, nonetheless excluding this information from quantitative evaluations.
@@ 2.2 Constructing a Dataset for PMD
Figure 2: Selected examples from Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD). For each matching pair of a video frame and procedural text (in this case, “Fold the cloth with your hands”), we generate a success example, and generate various types of mistake examples by sampling alternate video frames: incomplete execution of the procedure, execution with the wrong verb (e.g., wringing a cloth instead of folding), execution with the wrong noun (e.g., folding a paper instead of a cloth), and execution with both the wrong verb and noun (e.g., opening a notepad instead of folding a cloth). Images slightly cropped for space.
@@ 3 Evaluating Coherence in PMD
Next, we describe our application of a fine-tuned NLI model to calculate two coherence metrics for explainable PMD: relevance of questions and informativeness of answers to those questions.
@@
3.1 Using NLI Models to Judge Success
As visualized in Figure 3, LMs fine-tuned for NLI can estimate the sufficiency of visual questions and answers in explaining whether a procedure was successfully completed. To implement this, we need an NLI model $f_{e}$ (which returns a probability that an input premise string entails a hypothesis string), a premise transformation $t_{p}$ (which converts a question $Q$ and answer $A$ into a declarative statement to add to the premise), and a hypothesis prompt template $t_{h}$ (which creates a hypothesis about the success of the procedure in $T$). We can then calculate the NLI model’s probability for the success of procedure $T$ based on the explanation $E=(\mathcal{Q},\mathcal{A})$ as follows:
We implement $f_{e}$ with BART Lewis et al. (2020) fine-tuned on the large-scale MultiNLI dataset Williams et al. (2017), [^3] Model weights can be downloaded at https://huggingface.co/facebook/bart-large-mnli. applying softmax over its logits for entailment and contradiction to get an entailment probability. We follow Srinivasan et al. (2024) and prompt a foundational LM to implement $t_{p}$. [^4] To conserve GPU memory, we later use the evaluated VLM’s LM backbone for rephrasing. Details in Appendix B. For the procedural text $T$, we choose a success prompt template $t_{h}$ “The procedure $\langle P\rangle$ has been successfully executed.” [^5] While we could also use an LM to rephrase this statement, we found that more complex procedural texts were less likely to be rephrased accurately, and ultimately degraded the correlation with human judgments (see Appendix C).
@@ 3.2 Relevance
For evaluation across a dataset, we summarize the relevance of a sequence of questions generated for a particular example of explainable PMD by taking the mean relevance of generated questions with respect to previous questions and answers:
@@ 3.3 Informativeness
Since a relevant question does not guarantee an informative answer, [^7] For example, in the procedure “In a bowl, add the cut cherry tomatoes,” “Are there tomatoes in the bowl?” is a relevant question, but a “Yes” answer does not give us enough information to confirm that the procedure is 100% complete (e.g., there could be more tomatoes still outside the bowl). and VLM errors in answering questions could unintentionally introduce conflicting information, it is necessary to evaluate the sufficiency of VLMs’ answers in justifying a PMD decision. To achieve this, we measure the informativeness of a predicted answer $A^{\prime}$ for a question $Q^{\prime}$ to the success of a procedure $T$ (given previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$) as follows:
@@
$H$ is the binary Shannon entropy of the success probability $p_{e}$, calculated by $H(p)=-p\log_{2}p-(1-p)\log_{2}(1-p),p\in[0,1]$. This definition for informativeness quantifies how much information an answer to a question gives us in determining the success of the procedure. As such, if the success probability given this answer $A^{\prime}$ to $Q^{\prime}$ is confident (i.e., very low or high), this indicates that $A^{\prime}$ (along with $\mathcal{Q}$ and $\mathcal{A}$) are sufficient to make a decision, and thus informativeness will be high. On the other hand, a success probability closer to 50% suggests that the information gathered thus far is insufficient, yielding low informativeness. We express informativeness as a percentage (i.e., of 1, the maximum absolute value of bits here).
@@
We also wish to account for cases where the information gathered in questions and answers indicates the wrong PMD decision. To do so, we define the NLI model’s PMD belief $y_{e}(T|\mathcal{Q},\mathcal{A})$ as 1 (mistake) if $p_{e}(T|\mathcal{Q},\mathcal{A})<0.5$, else 0 (success). We then define the reference-adjusted informativeness:
@@
This makes informativeness negative if the NLI model judges the evidence gathered as indicating the wrong final answer for PMD.
@@
For evaluation across a dataset, we summarize the (reference-adjusted) informativeness of a set of questions and answers generated for a particular example by taking the maximum value across the self-dialog:
@@
4 Experimental Results
In this section, we present a broad set of experiments with VLMs in explainable PMD to validate our proposed metrics. First, we provide details for how VLMs are applied to the problem. We then introduce two inference-time interventions to encourage VLMs to select more coherent questions from generated candidates, as well as an approach to encourage them to generate more coherent candidates through preference optimization from our proposed coherence metrics. After evaluating and comparing these approaches, we show that our proposed metrics visualize and characterize common behaviors of VLMs, enabling a panoptic understanding of their capabilities in PMD.
4.1 Applying VLMs to Explainable PMD
Our experiments evaluate explainable PMD in InstructBLIP Dai et al. (2023a), LLaVA 1.5-7B Liu et al. (2023), and Llama 3.2-Vision-11B Dubey et al. (2024). [^8] See https://huggingface.co/Salesforce/instructblip-vicuna-7b, https://huggingface.co/liuhaotian/llava-v1.5-7b, and https://huggingface.co/meta-llama/Llama-3.2-11B-Vision. We intentionally target smaller, open-source VLMs that are feasible for online task guidance. The models apply different architectures, training datasets, and/or training strategies to integrate vision into the model. While LLaVA and InstructBLIP were not trained on Ego4D data, Dubey et al. (2024) does not reveal whether Llama 3 was. Details and prompt templates can be found in Appendix D, while key implementation details about the iterative self-dialog are provided below.
VQG.
We first prompt the VLM to generate a series of questions given the procedural text and the dialog history with previous questions and answers (enabling deductive reasoning). To encourage logical but diverse questions, we apply greedy beam search decoding to generate 4 candidates, from which we select the most likely candidate that has not already been generated.
VQA.
After a question is generated, it is answered by the VLM given only the question and video frame. If the probability of the chosen answer (i.e., “Yes” or “No”) exceeds an answer sureness threshold of 60%, we append it to the dialog history, otherwise we append “Unsure.” These unsure answers are excluded from the sets of questions $\mathcal{Q}$ and answers $\mathcal{A}$ in the calculation of example-level relevance and informativeness (defined in Section 3). Effectively, they are assigned zero informativeness and excluded from previous questions and answers in metric calculations.
Success classification.
Stopping criteria.
To prevent over-generating noisy information, which can degrade PMD accuracy, we implement an early stopping mechanism to determine whether to stop generating questions based on the success likelihood after each iteration. The self-dialog stops early (i.e., before $n$ questions have been generated) if the likelihood of success either stabilizes (i.e., changes by less than $\delta$ for two consecutive iterations) or becomes highly confident (i.e., subceeds $\epsilon$ or exceeds $1-\epsilon$). The hyperparameters $n$, $\delta$, and $\epsilon$ are tuned based on the validation data for each presented approach.
@@ 4.2 Coherent Question Selection
While a typical beam search would rank candidate questions by their likelihood, an alternative approach is to re-rank candidates using the reference-free coherence metrics introduced in Section 3. This could encourage the selection of questions that are most likely to bring in new, salient, and helpful information. Furthermore, as prior work has shown how the in-context learning capability of LMs enables them to accurately predict physical states from just a few in-context demonstrations Zhang et al. (2023), we will examine how supplementing candidate questions generated from the dialog history with additional questions generated through in-context learning approach impacts VLM performance.

### v2

# Transparent and Coherent Procedural Mistake Detection

@@ Abstract
Procedural mistake detection (PMD) is a challenging problem of classifying whether a human user (observed through egocentric video) has successfully executed a task (specified by a procedural text). Despite significant recent efforts, machine performance in the wild remains nonviable, and the reasoning processes underlying this performance are opaque. As such, we extend PMD to require generating visual self-dialog rationales to inform decisions. Given the impressive, mature image understanding capabilities observed in recent vision-and-language models (VLMs), we curate a suitable benchmark dataset for PMD based on individual frames. As our reformulation enables unprecedented transparency, we leverage a natural language inference (NLI) model to formulate two automated metrics for the coherence of generated rationales. We establish baselines for this reframed task, showing that while VLMs struggle off-the-shelf, their accuracy, coherence, and efficiency can be improved by incorporating these metrics into common inference and fine-tuning methods- though not without tradeoff. Lastly, our multi-faceted metrics visualize common outcomes, highlighting areas for further improvement.
Transparent and Coherent Procedural Mistake Detection
@@
Figure 1: To reason through the complex task of procedural mistake detection (PMD), vision-and-language models (VLMs) are conditioned to gather visual evidence through an iterative self-dialog to rationalize their final decision.
@@ 1 Introduction
The problem of interactive task guidance has recently attracted attention in the AI research community Bao et al. (2023); Wang et al. (2023a); Peddi et al. (2024); Bohus et al. (2024), stemming from significant efforts to build and learn from large-scale procedural video datasets Zhou et al. (2018); Damen et al. (2018); Miech et al. (2019); Grauman et al. (2022). A successful task guidance agent can observe a human user through video and interact with them through language to guide them to complete a task. A key component of such an agent is procedural mistake detection (PMD): the ability to detect when the user’s actions deviate from a procedural text, e.g., a recipe or instruction manual. To achieve this, a system must anticipate the success conditions for actions in this procedural text, then verify them by extracting relevant physical state information from the visual scene.
@@
Prior work in PMD has explored a mix of specialized (primarily vision-based) classifiers Sener et al. (2022); Wang et al. (2023a); Peddi et al. (2024) as well as foundational language models (LMs) and vision-and-language models (VLMs) Du et al. (2023); Bao et al. (2023); Peddi et al. (2024); Flaborea et al. (2024), but this problem has proven difficult, and current approaches fail to achieve accuracy above chance in zero-shot PMD in the wild. Qualitatively, Bao et al. (2023) find that while the web-scale multimodal pre-training of foundational VLMs enables coverage of a wide variety of procedures, they often produce noisy, vague, or otherwise insufficient information from visual scenes to facilitate PMD. This capability to extract and reason over key task-relevant visual information is crucial to PMD, but prior work has largely overlooked it, instead targeting binary or categorical classification tasks in their system design and quantitative evaluations. Consequently, the reasons for VLMs’ decisions are opaque, hindering practical use [^1] Since VLMs are unreliable mistake detectors, transparency is crucial for the user to understand the system’s reasoning process and act on or disregard it accordingly. and continued improvement.
@@
To promote the development of PMD systems with transparent and justified decisions, we propose a reformulated problem of coherent PMD: given a procedural text and egocentric video frame, VLMs must not only classify whether a mistake has occurred, but also support this decision with a rationale consisting of evidence from the visual scene. As shown in Figure 1, this rationale takes the form of an iterative self-dialog of generated questions and yes-no answers which directly condition classification. [^2] This choice of rationale resembles visual question answering (VQA; Antol et al., 2015) and visual dialog Das et al. (2017), long-studied multimodal tasks that VLMs excel at Dai et al. (2023a); Liu et al. (2023); Dubey et al. (2024). Since recent VLMs struggle to extract detailed information from videos, but have exhibited more mature image understanding capabilities, we curate an approachable large-scale dataset for PMD based on individual video frames annotated in Ego4D Grauman et al. (2022). We define two metrics for the coherence of generated rationales based on a natural language inference (NLI) model. To lay a foundation for research in coherent PMD, we establish baselines by exploring three natural interventions to VLMs: (1) we use our metrics to re-rank candidate questions generated by VLMs, (2) we harness VLMs’ in-context learning capability to generate additional candidate questions based on human-written examples, and (3) we use our metrics to fine-tune VLMs to generate more coherent questions. Our results show that while VLMs struggle off-the-shelf, these interventions can improve VLMs’ accuracy, coherence, and rationale generation efficiency, albeit creating tradeoffs between these properties. We lastly show how our multi-faceted metrics visualize common outcomes in coherent PMD, enabling fine-grained evaluation and identification of areas for future improvement.
Figure 2: Selected examples from Ego4D Grauman et al. (2022) for Procedural Mistake Detection (Ego4D-PMD). For each matching pair of a video frame and procedural text, we generate a success example and various mistake examples by sampling alternate video frames: incomplete execution, execution with the wrong verb (e.g., wringing a cloth instead of folding), execution with the wrong noun (e.g., folding paper instead of a cloth), and execution with both the wrong verb and noun (e.g., opening a notepad instead of folding a cloth). Images cropped for space.
@@ 2 Problem Formulation and Dataset
In this section, we define the extended problem of coherent PMD in an approachable manner for VLMs, describe how to apply VLMs to the problem, then lastly introduce a benchmark dataset we curated for evaluating coherent PMD.
2.1 Defining Coherent PMD
The inputs for PMD are a short procedural text $T$ and a single video frame $F$, which may or may not visualize the successful completion of the procedure described in $T$. Given these inputs, a system should return a binary success decision $y$ for whether the procedure has been successfully completed ($y=0$ indicates success, and $y=1$ indicates the detection of a mistake). In coherent PMD, it must additionally generate a rationale $R=(\mathcal{Q},\mathcal{A})$, where $\mathcal{Q}=\{Q_{1},Q_{2},\dots,Q_{n}\}$ and $\mathcal{A}=\{A_{1},A_{2},\dots,A_{n}\}$ are respectively sequences of $n$ yes-no questions and their predicted answers, which provide evidence for the decision.
2.2 Applying VLMs to Coherent PMD
As shown in Figure 1, we elicit a rationale from VLMs through a self-generated visual dialog Das et al. (2017) consisting of visual question generation (VQG) based on the procedural text and dialog history, and visual question answering (VQA) based on the video frame. The evidence compiled in this rationale then conditions a success classification for whether the procedure has been completed in the given video frame. This structure goes beyond past approaches for PMD using VLMs; while Du et al. (2023) only elicited success classification, Bao et al. (2023) used procedure-agnostic prompts to caption images before classification, nonetheless disregarding this information in quantitative evaluations.
VQG.
We prompt the VLM to generate a series of questions given the procedural text (and previous questions and answers in later iterations, enabling deductive reasoning). To encourage logical, diverse questions, we apply greedy beam search to generate 4 candidates, from which we select the most likely candidate not generated previously.
VQA.
After a question is generated, it is answered by the VLM given only the question and video frame. If the probability of the chosen answer (i.e., “Yes” or “No”) exceeds an answer sureness threshold of 60%, we append it to the dialog history, otherwise we append “Unsure.” [^3] Unsure answers are excluded from example-level informativeness (defined in Section 3), and excluded from previous questions and answers in metric calculations.
Success classification.
Stopping criteria.
To prevent over-generating evidence, which could introduce noise and degrade PMD accuracy, we implement an early stopping mechanism to determine whether to stop generating questions based on the success likelihood after each iteration. The self-dialog stops early (i.e., before $n$ questions are generated) if the success likelihood stabilizes (i.e., changes by less than $\delta$ for two consecutive iterations) or becomes highly confident (i.e., subceeds $\epsilon$ or exceeds $1-\epsilon$). $n$ is fixed at 10, while $\delta$ and $\epsilon$ are tuned based on the validation data for each presented approach.
@@ 3 Evaluating Coherence in PMD
Next, we describe our application of a fine-tuned NLI model to calculate two evaluation metrics for coherent PMD: relevance of questions and informativeness of answers to those questions.
3.1 Using NLI Models to Judge Success
As shown in Figure 3, LMs fine-tuned for NLI can estimate the sufficiency of visual questions and answers in rationalizing whether a procedure was successfully completed. This requires an NLI model $f_{e}$ (which returns a probability that an input premise string entails a hypothesis string), a premise transformation $t_{p}$ (which converts a question $Q$ and answer $A$ into a declarative statement to add to the premise), and a hypothesis prompt template $t_{h}$ (which creates a hypothesis about the success of the procedure in $T$). We then calculate the NLI model’s probability for the success of procedure $T$ based on the rationale $R=(\mathcal{Q},\mathcal{A})$:
We implement $f_{e}$ with BART Lewis et al. (2020) fine-tuned on the large-scale MultiNLI dataset Williams et al. (2017), [^4] See https://huggingface.co/facebook/bart-large-mnli. applying softmax over its logits for entailment and contradiction to get an entailment probability. We follow Srinivasan et al. (2024) and prompt a foundational LM to implement $t_{p}$. [^5] To conserve GPU memory, we later use the evaluated VLM’s LM backbone for rephrasing. Details in Appendix B. For the procedural text $T$, we choose a success prompt template $t_{h}$ “The procedure $\langle P\rangle$ has been successfully executed.” [^6] A template was used as we found complex procedural texts were unlikely to be rephrased accurately, degrading correlation with human judgments (see Appendix C).
@@ 3.2 Relevance
To reward systems that propose consistently relevant questions in our evaluations, we summarize the relevance of a sequence of questions generated for a particular example by taking the mean question relevance with respect to previous questions and answers:
@@ 3.3 Informativeness
Since a relevant question does not guarantee an informative answer, [^8] For example, in the procedure “In a bowl, add the cut cherry tomatoes,” “Are there tomatoes in the bowl?” is a relevant question, but a “Yes” answer is insufficient to confirm 100% completion (more tomatoes could be outside the bowl). and VLM errors in answering questions could unintentionally introduce conflicting information, it is necessary to evaluate the sufficiency of VLMs’ answers in justifying a PMD decision. To achieve this, we measure the informativeness of a predicted answer $A^{\prime}$ for a question $Q^{\prime}$ to the success of a procedure $T$ (given previous questions and answers $\mathcal{Q}$ and $\mathcal{A}$) as follows:
@@
$H$ is the binary Shannon entropy of the success probability $p_{e}$, calculated by $H(p)=-p\log_{2}p-(1-p)\log_{2}(1-p),p\in[0,1]$. This definition for informativeness quantifies how much information an answer to a question provides in determining the success of the procedure. As such, if the success probability given this answer $A^{\prime}$ to $Q^{\prime}$ is confident (i.e., very low or high), this indicates that $A^{\prime}$ (along with $\mathcal{Q}$ and $\mathcal{A}$) are sufficient to make a decision, and thus informativeness is high. Conversely, a success probability close to 50% suggests the evidence gathered thus far is insufficient, yielding low informativeness. Informativeness is expressed as a number of bits between 0 and 1.
@@
We also wish to account for cases where the evidence gathered in questions and answers indicates the wrong PMD decision. To do so, we define the NLI model’s PMD belief $y_{e}(T|\mathcal{Q},\mathcal{A})$ as 1 (mistake) if $p_{e}(T|\mathcal{Q},\mathcal{A})<0.5$, else 0 (success). Given the ground truth PMD label $y^{*}$, we then define the reference-adjusted informativeness to be negative if the NLI model judges the evidence gathered as indicating the wrong success decision:
@@
To summarize the sufficiency of information gathered throughout the self-dialog (which may have some uninformative answers), we take the maximum (reference-adjusted) informativeness across the self-dialog:
@@
4 Rationale Coherence Interventions
To validate these metrics and examine the relationship between rationale coherence, PMD accuracy, and rationale generation efficiency, we next introduce two inference-time interventions to encourage VLMs to select more coherent questions from generated candidates, as well as a preference optimization approach to encourage generating more coherent candidates based on coherence metrics.
@@ 4.1 Coherent Question Selection
While a typical beam search would use a sequence likelihood-based approach to rank candidate questions, an alternative is to re-rank candidates using the reference-free coherence metrics introduced in Section 3. This could encourage selecting questions that are likely to bring in new, salient, and helpful information. Furthermore, as adaptive information-seeking is a core component of humans’ reasoning capabilities Coenen et al. (2019), we will supplement candidate questions with candidates generated through in-context learning from human-written examples (which we assume are reasonably coherent and effective).